The sum of the largest and smallest of three consecutive even numbers is 68. What are the three numbers?
Answer by Shana-D77(122)

To generalize three consecutive even numbers, we can write:
first number: x
second number number: x + 2
third number: x + 4
So, the largest is x + 4. The smallest is x. "the sum of the largest and smallest is 68":
x + (x + 4) = 68
2x + 4 = 68 (combined like terms on the left side of the equation)
2x = 64 (subtracted 4 from both sides)
x = 32 (divided bith sides by 2)
If the smallest number (x) is 32, then the next even number (x + 2) is 34, and the next one after that (x + 4) is 36.
check: 32 + 36 = 68
answer: 32, 34, 36